OpenStudy (anonymous):

A ballon has a circumfrence of 11cm. Use the circumference to approximate the surface are of the balloon to the nearest square centimeter! Help! a 121 b380 c39 d154

OpenStudy (anonymous):

use this concept http://answers.yahoo.com/question/index?qid=20110715171126AAXwqEF

OpenStudy (anonymous):

I looked at that I couldn't figure it out.

OpenStudy (anonymous):

just plug your info to the equation that the person gave

OpenStudy (anonymous):

Surface area is 4 pi r sq Circumference is 2 pi r Circumference is 23 cm 2pi r =23 r = 23 x 7/44 = 3.659 cm Surface area = 4 pi r sq = 4 x 22/7 x 3.659 x 3.659 = 168.31 Answer : 168 sq cm.

OpenStudy (anonymous):

where is the 7 and 44 coming from though? that is what mixed me up

HanAkoSolo (jamierox4ev3r):

is this ballon a perfect sphere?

HanAkoSolo (jamierox4ev3r):

@zacmcgraw8 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

oh okay then the surface area of a sphere is \[A=4\pi(r)^2\]

HanAkoSolo (jamierox4ev3r):

divide the diameter by 2 and you have the radius. Then plug it in to the equation

HanAkoSolo (jamierox4ev3r):

A=4pi(5.5)^2 A=4pi(30.25) A=4 (94.985 A=379.94

HanAkoSolo (jamierox4ev3r):

if you round this to the nearest square acre, then your answer is 380 acres ^2. This means your answer is B.

HanAkoSolo (jamierox4ev3r):

Does this make sense @zacmcgraw8 ?

OpenStudy (anonymous):

Yes